OpenStudy (anonymous):

If f(x) = x + 2 and g(x) = 3x, find f of x divided by g of x

OpenStudy (anonymous):

\[\frac{ x+2 }{ 3x }\]

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }x+\frac{1}{3}\]